OpenStudy (anonymous):

PLEASE HELP MEDALLLLLLLLLLLLLLLLLLll What is the slope of the line that passes through the points (14, 5) and (20, –4)? A. -3/2 B. -2/3 C. 2/3 D. 3/2

OpenStudy (black_ninja123):

Do you know the slope formula

OpenStudy (tabdo1226):

are you an anonyumus hacker me too

OpenStudy (black_ninja123):

\[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (black_ninja123):

plug in the numbers and let me know what you get

OpenStudy (anonymous):

SO...

OpenStudy (anonymous):

23.5?

OpenStudy (black_ninja123):

one sec

OpenStudy (black_ninja123):

thats wrong

OpenStudy (black_ninja123):

It showed by -1.5

OpenStudy (black_ninja123):

should*

OpenStudy (anonymous):

thats the only answer i got tho

OpenStudy (black_ninja123):

nvm its -3/2

OpenStudy (anonymous):

ok thanks man

OpenStudy (anonymous):

100% ;)

OpenStudy (black_ninja123):

Np

OpenStudy (black_ninja123):

efrdftrfdfrtfgvrdf4erd

OpenStudy (tabdo1226):

http://openstudy.com/study#/updates/5554d015e4b0db8ecf8e7515 need help

OpenStudy (astrophysics):

\[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} } \implies \frac{ -4-5 }{ 20-14 } \implies \frac{ -9 }{ 6 } = \frac{ -3 }{ 2 }\]

OpenStudy (astrophysics):

Just to clear it up